# Comprehensive Review of Neural Differential Equations for Time Series Analysis

### v1

@@ 5.2 NCDE Methods
Neural CDE (Kidger et al., 2020) models the latent state $\bm{z}(t)$ with a piecewise-smooth control path $X(t)$. The Riemann–Stieltjes integral allows Neural CDEs to handle irregular or asynchronous time series data effectively, making them suitable for tasks like interpolation and forecasting. Neural Rough Differential Equations (Neural RDEs) (Morrill et al., 2021a) extend Neural CDEs by using the log-signature of the control path $\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)$ with $X:[t_{0}:t_{n}]$ and certain interval $t_{0}\leq r_{i}<r_{i+1}\leq t_{n}$, capturing higher-order variations. This enables Neural RDEs to handle rough paths, improving robustness for long and complex time series. Recently, Walker et al. (2024) further extended rough path theory to Log-NCDEs. These approaches leverage continuous-time dynamics, enabling flexible modeling of time series.
Recent works (Jhin et al., 2022, 2023) emphasize training-based approaches for constructing control paths in NCDEs. For instance, Attentive Neural Controlled Differential Equation (ANCDE) (Jhin et al., 2024) extend this line of research by incorporating attention mechanisms for dynamic path construction. This is implemented using two coupled NCDEs: one computes attention values, $\bm{a}(t)=\sigma(\bm{h}(t))$ using a activation $\sigma$, and the other applies the attention-modulated control path, $Y(t)$, which is combination of vector $a(t)$ and $X(t)$.
DualDynamics  (Oh et al., 2025) combines explicit and implicit mechanisms within a unified framework to enhance the modeling of irregular time series data. By integrating Neural ODEs for explicit time evolution and learnable implicit updates for latent state transitions, DualDynamics achieves a balance between interpretability and robustness. These variations of NCDEs have been extensively applied to irregularly-sampled time series or data with missing observations, demonstrating their effectiveness in variety of tasks including classification, imputation, and forecasting.
@@ 5.3 NSDE Methods
Neural SDEs, parameterized by neural networks for both drift and diffusion terms, provide a flexible framework for modeling complex stochastic processes, with foundational contributions from Han et al. (2017) and Tzen and Raginsky (2019). Neural Jump SDE (Jia and Benson, 2019) incorporates Poisson jump processes to model discontinuous state changes. Zhang et al. (2024) extended this to temporal point processes. While requiring specialized solvers, this approach effectively captures rare, discontinuous events in NSDEs.
@@
Latent SDE (Li et al., 2020) extends the encoding-decoding paradigm of Latent ODE (Rubanova et al., 2019) by incorporating stochastic dynamics, allowing for more flexible modeling of uncertainty and variability. Latent SDE defines a prior process $\widetilde{\bm{z}}(t)$ and an approximate posterior process $\bm{z}(t)$, both parameterized as Neural SDEs with $h_{\theta}$, $h_{\phi}$, and $\sigma$. During the training, the Kullback–Leibler (KL) divergence measures and minimizes the prior-posterior discrepancy while maintaining uncertainty. On the other hand, Kidger et al. (2021b) extends Wasserstein generative adversarial network (GAN) to continuous-time stochastic processes, where both generator and discriminator are formulated as Neural SDEs.
@@
Oh et al. (2024) introduced Stable Neural SDEs with three variants: Neural Langevin-type SDE (LSDE) with additive noise for invariant measures, Neural Linear Noise SDE (LNSDE) with multiplicative noise for scale-dependent fluctuations, and Neural Geometric SDE (GSDE) with state-dependent noise for exponential dynamics. Note that, these formulations utilize the augmented states with control path, $\overline{{\bm{z}}}(t)=\zeta(t,{\bm{z}}(t),X(t);\theta_{\zeta})$, in order to ensure stability for irregular time series analysis. Therefore, NSDEs extend NODEs by incorporating stochasticity, enabling robust modeling of complex, uncertain, and discontinuous dynamics.
@@ 6 Discussion and Future Directions
NDEs face scalability challenges with continuous-time solvers. Fixed integrators or adaptive integrators, while accurate, can lead to unpredictable computation times in stiff regions (Chen et al., 2018; Kim et al., 2021; Kidger, 2021). Specialized techniques like parallel-in-time integration and GPU acceleration offer promising solutions (Rackauckas et al., 2020; Gholami et al., 2019). Future research should focus on developing comprehensive integration schemes that can automatically balance computational efficiency with numerical accuracy, particularly for large-scale applications in real-time settings.
@@
Current theoretical understanding requires expansion, particularly for non-stationary and noisy data. While stability guarantees exist for NODEs and NCDEs through Lipschitz constraints (Massaroli et al., 2020b), similar guarantees for NSDEs remain incomplete (Oh et al., 2024). Drift-diffusion analysis under strong noise conditions needs further investigation (Tzen and Raginsky, 2019; Li et al., 2020). Development of rigorous frameworks for analyzing convergence properties and error bounds in these stochastic settings would significantly advance both theoretical foundations and practical applications of NDEs.
@@
Physics-Informed NDEs (or Physics Informed Neural Networks) incorporate domain knowledge, such as Partial Differential Equations (PDEs) constraints and conservation laws, to align modeled dynamics with physical systems. Recent advances in physics-informed architectures have demonstrated remarkable success in capturing multi-scale phenomena and handling noisy measurements in complex physical systems (Rudy et al., 2017; Raissi et al., 2019; Cuomo et al., 2022). By embedding structural priors, these approaches ensure interpretability and reliability in real-world tasks.
@@
Hybrid architectures combine NDEs with models like graph neural networks or Transformers to capture complex time series patterns. Graph-based NDEs handle spatial-temporal data (Poli et al., 2019; Choi et al., 2022), while attention mechanism address long-range dependencies and irregular sampling (Jhin et al., 2024; Li et al., 2024). Furthermore, Oh et al. (2025) integrates explicit and implicit mechanisms for irregular time series analysis. Future research directions should focus on developing more efficient training algorithms for these hybrid architectures, investigating theoretical guarantees, and exploring applications in real-world domains where both temporal dynamics and structural relationships play crucial roles.
@@ 7 Conclusion
NDEs represent a significant advancement in time series modeling, offering a principled approach to handling continuous dynamics. Through various formulations—NODEs, NCDEs, and NSDEs—they provide flexible frameworks for complex temporal data, including cases with irregular sampling or missing data. While challenges remain in computational efficiency and theoretical understanding, ongoing developments in solver techniques, stability analysis, and hybrid architectures continue to enhance their capabilities. Beyond current applications in classification, interpolation and forecasting, NDEs show potential in anomaly detection, reinforcement learning, and multi-agent systems. As the field evolves, NDEs are increasingly positioned to address sophisticated time series challenges across diverse applications.

### v2

@@ 5.2 NCDE Methods
Neural CDE Kidger et al. (2020) models the latent state $\bm{z}(t)$ with a piecewise-smooth control path $X(t)$. The Riemann–Stieltjes integral allows Neural CDEs to handle irregular or asynchronous time series data effectively, making them suitable for tasks like interpolation and forecasting. Neural Rough Differential Equations (Neural RDEs) Morrill et al. (2021a) extend Neural CDEs by using the log-signature of the control path $\mathrm{LogSig}_{[r_{i},r_{i+1}]}(X)$ with $X:[t_{0}:t_{n}]$ and certain intervals $t_{0}\leq r_{i}<r_{i+1}\leq t_{n}$, capturing higher-order variations. This enables Neural RDEs to handle rough paths, improving robustness for long and complex time series. Recently, Walker et al. (2024) further extended rough path theory to Log-NCDEs. These approaches leverage continuous-time dynamics, enabling flexible modeling of time series.
Recent works Jhin et al. (2022, 2023) emphasize training-based approaches for constructing control paths in NCDEs. For instance, Attentive Neural Controlled Differential Equation (ANCDE) Jhin et al. (2024) extend this line of research by incorporating attention mechanisms for dynamic path construction. This is implemented using two coupled NCDEs: one computes attention values, $\bm{a}(t)=\sigma(\bm{h}(t))$ using an activation $\sigma$, and the other applies the attention-modulated control path, $Y(t)$, which is a combination of vector $a(t)$ and $X(t)$. Similarly, Li et al. (2023) propose Neural LAD (Latent dynamics), which incorporates an attention-based transformation of the input path gradient $g_{v}\tfrac{dX}{dt}$ for continuous-time latent dynamics. In addition, Neural LAD employs a seasonality-trend transformation $h_{w}{(t)}$ to capture periodic and slowly varying patterns, and uses $\phi_{\gamma}\bigl{(}\bm{z}(t)\bigr{)}$ to learn structural relationships within the latent state.
DualDynamics  Oh et al. (2025) combines explicit and implicit mechanisms within a unified framework to improve the modeling of irregular time series data. By integrating Neural ODEs for explicit time evolution and learnable implicit updates for latent state transitions, DualDynamics achieves a balance between interpretability and robustness. These variations of NCDEs have been extensively applied to irregularly-sampled time series or data with missing observations, demonstrating their effectiveness in a variety of tasks including classification, imputation, and forecasting.
@@ 5.3 NSDE Methods
Neural SDEs, parameterized by neural networks for both drift and diffusion terms, provide a flexible framework for modeling complex stochastic processes, with foundational contributions from Han et al. (2017) and Tzen and Raginsky (2019). Neural Jump SDE Jia and Benson (2019) incorporates Poisson jump processes to model discontinuous state changes. Zhang et al. (2024) extended this to temporal point processes. While requiring specialized solvers, this approach effectively captures rare, discontinuous events in NSDEs.
@@
Latent SDE Li et al. (2020) extends the encoding-decoding paradigm of Latent ODE Rubanova et al. (2019) by incorporating stochastic dynamics, allowing for more flexible modeling of uncertainty and variability. Latent SDE defines a prior process $\widetilde{\bm{z}}(t)$ and an approximate posterior process $\bm{z}(t)$, both parameterized as Neural SDEs with $h_{\theta}$, $h_{\phi}$, and $\sigma$. During the training, the Kullback–Leibler (KL) divergence measures and minimizes the prior-posterior discrepancy while maintaining uncertainty. On the other hand, Kidger et al. (2021b) extends Wasserstein generative adversarial network (GAN) to continuous-time stochastic processes, where both generator and discriminator are formulated as Neural SDEs.
@@
Stable Neural SDEs, introduced in Oh et al. (2024), extend Neural SDEs to address stability and robustness in complex systems: Neural Langevin-type SDE (LSDE) with additive noise for invariant measures, Neural Linear Noise SDE (LNSDE) with multiplicative noise for scale-dependent fluctuations, and Neural Geometric SDE (GSDE) with state-dependent noise for exponential dynamics. Note that, these formulations utilize the augmented states with control path, $\overline{{\bm{z}}}(t)=\zeta(t,{\bm{z}}(t),X(t);\theta_{\zeta})$, in order to ensure stability for irregular time series analysis. Therefore, NSDEs extend NODEs by incorporating stochasticity, enabling robust modeling of complex, uncertain, and discontinuous dynamics.
5.4 Empirical Comparison
We present a consolidated summary of diverse empirical results from Oh et al. (2024) and Oh et al. (2025), encompassing tasks such as interpolation on PhysioNet Mortality, forecasting on MuJoCo, and classification on PhysioNet Sepsis.
We summarize interpolation experiments on the 2012 PhysioNet Mortality dataset Silva et al. (2012), which contains 37 ICU variables recorded over 48 hours. Following Rubanova et al. (2019), timestamps were rounded to the nearest minute, resulting in up to 2880 intervals. As in Shukla and Marlin (2021), observation rates varied from 50% to 90%, using 8000 instances to predict the unobserved values. Benchmark models included RNN-VAE Chen et al. (2018), mTAND-Full Shukla and Marlin (2021), L-ODE-RNN Chen et al. (2018), L-ODE-ODE Rubanova et al. (2019), DualDynamics Oh et al. (2025), Latent SDE Li et al. (2020), and variants of Neural SDEs Oh et al. (2024). Well-designed NDE methods outperform conventional models in interpolation performance, as shown in Table 2.
Methods Observed %
50% 60% 70% 80% 90%
RNN-VAE 13.418 $\pm$0.008 12.594 $\pm$0.004 11.887 $\pm$0.005 11.133 $\pm$0.007 11.470 $\pm$0.006
mTAND-Full 4.139 $\pm$0.029 4.018 $\pm$0.048 4.157 $\pm$0.053 4.410 $\pm$0.149 4.798 $\pm$0.036
L-ODE-RNN 8.132 $\pm$0.020 8.140 $\pm$0.018 8.171 $\pm$0.030 8.143 $\pm$0.025 8.402 $\pm$0.022
L-ODE-ODE 6.721 $\pm$0.109 6.816 $\pm$0.045 6.798 $\pm$0.143 6.850 $\pm$0.066 7.142 $\pm$0.066
DualDynamics 3.631 $\pm$0.049 3.659 $\pm$0.028 3.463 $\pm$0.032 3.224 $\pm$0.044 3.114 $\pm$0.050
Latent SDE 8.862 $\pm$0.036 8.864 $\pm$0.058 8.686 $\pm$0.122 8.716 $\pm$0.032 8.435 $\pm$0.077
Neural SDE 8.592 $\pm$0.055 8.591 $\pm$0.052 8.540 $\pm$0.051 8.318 $\pm$0.010 8.252 $\pm$0.023
Neural LSDE 3.799 $\pm$0.055 3.584 $\pm$0.055 3.457 $\pm$0.078 3.262 $\pm$0.032 3.111 $\pm$0.076
Neural LNSDE 3.808 $\pm$0.078 3.617 $\pm$0.129 3.405 $\pm$0.089 3.269 $\pm$0.057 3.154 $\pm$0.084
Neural GSDE 3.824 $\pm$0.088 3.667 $\pm$0.079 3.493 $\pm$0.024 3.287 $\pm$0.070 3.118 $\pm$0.065
Table 2: Interpolation performance (Test MSE ×$\times 10^{-3}$) versus percent observed time points on PhysioNet Mortality dataset
For forecasting and classification task, we consider a variety of NDE-related methods, including GRU-$\Delta t$ Choi et al. (2016), GRU-D Che et al. (2018), GRU-ODE De Brouwer et al. (2019), ODE-RNN Rubanova et al. (2019), Latent-ODE Rubanova et al. (2019), Augmented-ODE Dupont et al. (2019), Attentive co-evolving neural ordinary differential equations (ACE-NODE) Jhin et al. (2021), Neural CDE Kidger et al. (2020), ANCDE Jhin et al. (2024), EXtrapolation and InTerpolation-based model (EXIT) Jhin et al. (2022), LEArnable Path-based model (LEAP) Jhin et al. (2023), DualDynamics Oh et al. (2025), and variants of Neural SDEs Oh et al. (2024) including Neural LSDE, Neural LNSDE, and Neural GSDE.
Table 3 summarizes forecasting results, which used the MuJoCo dataset Todorov et al. (2012) with the Hopper setup from the DeepMind control suite Tassa et al. (2018). Following Jhin et al. (2022, 2023), models were trained to forecast 10 future steps from 50 observed points under 30%, 50%, and 70% missingness. Similar to the previous results, NDE models demonstrate both superior prediction performance under varying missingness and greater memory efficiency in general, in MuJoCo forecasting.
Methods Test MSE Memory (MB)
0% dropped 30% dropped 50% dropped 70% dropped
GRU-$\Delta t$ 0.223 $\pm$0.020 0.198 $\pm$0.036 0.193 $\pm$0.015 0.196 $\pm$0.028 533
GRU-D 0.578 $\pm$0.042 0.608 $\pm$0.032 0.587 $\pm$0.039 0.579 $\pm$0.052 569
GRU-ODE 0.856 $\pm$0.016 0.857 $\pm$0.015 0.852 $\pm$0.015 0.861 $\pm$0.015 146
ODE-RNN 0.328 $\pm$0.225 0.274 $\pm$0.213 0.237 $\pm$0.110 0.267 $\pm$0.217 115
Latent-ODE 0.029 $\pm$0.011 0.056 $\pm$0.001 0.055 $\pm$0.004 0.058 $\pm$0.003 314
Augmented-ODE 0.055 $\pm$0.004 0.056 $\pm$0.004 0.057 $\pm$0.005 0.057 $\pm$0.005 286
ACE-NODE 0.039 $\pm$0.003 0.053 $\pm$0.007 0.053 $\pm$0.005 0.052 $\pm$0.006 423
Neural CDE 0.028 $\pm$0.002 0.027 $\pm$0.000 0.027 $\pm$0.001 0.026 $\pm$0.001 52.1
ANCDE 0.026 $\pm$0.001 0.025 $\pm$0.001 0.025 $\pm$0.001 0.024 $\pm$0.001 79.2
EXIT 0.026 $\pm$0.000 0.025 $\pm$0.004 0.026 $\pm$0.000 0.026 $\pm$0.001 127
LEAP 0.022 $\pm$0.002 0.022 $\pm$0.001 0.022 $\pm$0.002 0.022 $\pm$0.001 144
DualDynamics 0.006 $\pm$0.001 0.008 $\pm$0.001 0.008 $\pm$0.001 0.008 $\pm$0.001 318
Neural SDE 0.028 $\pm$0.004 0.029 $\pm$0.001 0.029 $\pm$0.001 0.027 $\pm$0.000 234
Neural LSDE 0.013 $\pm$0.000 0.014 $\pm$0.001 0.013 $\pm$0.000 0.013 $\pm$0.001 249
Neural LNSDE 0.012 $\pm$0.001 0.014 $\pm$0.001 0.013 $\pm$0.001 0.014 $\pm$0.000 273
Neural GSDE 0.013 $\pm$0.001 0.013 $\pm$0.001 0.013 $\pm$0.000 0.014 $\pm$0.000 306
Table 3: Forecasting MSE and memory usage versus percent dropped time points on MuJoCo dataset
Table 4 summarizes classification results, which used the PhysioNet Sepsis dataset Reyna et al. (2019), containing 40,335 patients and 34 temporal variables. Following Kidger et al. (2020), we compared classification with and without observation intensity (OI), an index reflecting patient severity. Classification performance was measured by AUROC. On the PhysioNet Sepsis task, NDE-based models consistently achieve higher AUROC with significantly lower memory usage compared to conventional baselines.
Methods Test AUROC Memory (MB)
OI No OI OI No OI
GRU-$\Delta t$ 0.878 $\pm$0.006 0.840 $\pm$0.007 837 826
GRU-D 0.871 $\pm$0.022 0.850 $\pm$0.013 889 878
GRU-ODE 0.852 $\pm$0.010 0.771 $\pm$0.024 454 273
ODE-RNN 0.874 $\pm$0.016 0.833 $\pm$0.020 696 686
Latent-ODE 0.787 $\pm$0.011 0.495 $\pm$0.002 133 126
ACE-NODE 0.804 $\pm$0.010 0.514 $\pm$0.003 194 218
Neural CDE 0.880 $\pm$0.006 0.776 $\pm$0.009 244 122
ANCDE 0.900 $\pm$0.002 0.823 $\pm$0.003 285 129
EXIT 0.913 $\pm$0.002 0.836 $\pm$0.003 257 127
DualDynamics 0.918 $\pm$0.003 0.873 $\pm$0.004 453 233
Neural SDE 0.799 $\pm$0.007 0.796 $\pm$0.006 368 240
Neural LSDE 0.909 $\pm$0.004 0.879 $\pm$0.008 373 436
Neural LNSDE 0.911 $\pm$0.002 0.881 $\pm$0.002 341 445
Neural GSDE 0.909 $\pm$0.001 0.884 $\pm$0.002 588 280
Table 4: AUROC and memory usage on PhysioNet Sepsis
Across diverse time series related tasks, NDE-based methods consistently achieve better performance compared to conventional models. Their advantages are particularly pronounced in dynamic prediction settings and irregularly sampled data, confirming their suitability for real-world time series modeling in a continuous-time manner.
@@ 6 Discussion and Future Directions
NDEs face scalability challenges with continuous-time solvers. Fixed-step integrators or adaptive integrators, while accurate, can lead to unpredictable computation times in stiff regions Chen et al. (2018); Kim et al. (2021); Kidger (2021). Specialized techniques like parallel-in-time integration and GPU acceleration offer promising solutions Rackauckas et al. (2020); Gholami et al. (2019). Future research should focus on developing comprehensive integration schemes that can automatically balance computational efficiency with numerical accuracy, particularly for large-scale applications in real-time settings.
@@
Current theoretical understanding requires expansion, particularly for non-stationary and noisy data. While stability guarantees exist for NODEs and NCDEs through Lipschitz constraints Massaroli et al. (2020b), similar guarantees for NSDEs remain incomplete Oh et al. (2024). Drift-diffusion analysis under strong noise conditions needs further investigation Tzen and Raginsky (2019); Li et al. (2020). Development of rigorous frameworks for analyzing convergence properties and error bounds in these stochastic settings would significantly advance both theoretical foundations and practical applications of NDEs.
@@
Physics-Informed NDEs (or Physics-Informed Neural Networks) incorporate domain knowledge, such as Partial Differential Equations (PDEs) constraints and conservation laws, to align modeled dynamics with physical systems. Recent advances in physics-informed architectures have demonstrated remarkable success in capturing multi-scale phenomena and handling noisy measurements in complex physical systems Rudy et al. (2017); Raissi et al. (2019); Cuomo et al. (2022). By embedding structural priors, these approaches ensure interpretability and reliability in real-world tasks.
@@
Hybrid architectures combine NDEs with models like graph neural networks or Transformers to capture complex time series patterns. Graph-based NDEs handle spatio-temporal data Poli et al. (2019); Choi et al. (2022), while attention mechanisms address long-range dependencies and irregular sampling Jhin et al. (2024); Li et al. (2024). Furthermore, Oh et al. (2025) integrates explicit and implicit mechanisms for irregular time series analysis. Future research directions should focus on developing more efficient training algorithms for these hybrid architectures, and exploring applications in real-world domains where both temporal dynamics and structural relationships are important.
@@ 7 Conclusion
This paper has provided a focused and comprehensive survey of NDEs, specifically for time series analysis. While existing literature includes reviews on broader topics or specific niches within this domain, a systematic consolidation and comparative analysis of the core architectural families and their variants, tailored to the unique challenges of temporal data, has been largely absent. Our work addresses this gap, offering a structured synthesis of advancements and outlining future research directions in this new paradigm.
NDEs represent a significant advancement in time series modeling, offering a principled approach to handling continuous dynamics. Through various formulations, including NODEs, NCDEs, and NSDEs, they provide flexible frameworks for complex temporal data, including cases with irregular sampling or missing data. While challenges remain in computational efficiency and theoretical understanding, ongoing developments in solver techniques, stability analysis, and hybrid architectures continue to enhance their capabilities. Beyond current applications in classification, interpolation, and forecasting, NDEs show potential in anomaly detection, reinforcement learning, and multi-agent systems. As the field evolves, NDEs are increasingly positioned to address sophisticated time series challenges across diverse applications.
Acknowledgments
[…]
This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00407852); the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health and Welfare, Republic of Korea (HI19C1095); the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00563597); the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828); and the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00443780, Development of Foundation Models for Bioelectrical Signal Data and Validation of Their Clinical Applications: A Noise-and-Variability Robust, Generalizable Self-Supervised Learning Approach).